The proportion of eligible voters in the next corporation election (in tumkur) who will vote for the incumbent is assumed to be 55%. What is the probability that in a random sample of 500 voters less than 49% say they will vote for the incumbent? Write your conclusion.


1
Expert's answer
2022-03-01T12:33:48-0500

"\hat{p}" is normally distributed with a mean of "\mu_{\hat{p}}=p" and a standard deviation

"\sigma_{\hat{p}}=\sqrt{\dfrac{p(1-p)}{n}}"

Given "p=0.55, n=500."


"P(\hat{p}<0.49)=P\bigg(Z<\dfrac{0.49-0.55}{\sqrt{\dfrac{0.55(1-0.55)}{500}} }\bigg)"

"\approx P(Z<-2.6968)\approx0.0035"